make y the subject.

x=2/3y^2+4
1
Expert's answer
2011-10-07T11:31:51-0400
Let's subtract 4 from each part of equation:
x-4 = 2/3y^2
Trransfering the multiplier:
3/2(x-4) = y^2
So, y = (3/2(x-4))^0.5